Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2022 November Paper 52 Q5

A Level / CIE / S1

CIE 9709 2022 Nov Paper 52 Paper · Question 5

题目

Problem

Eric has three coins. One of the coins is fair. The other two coins are each biased so that the probability of obtaining a head on any throw is 14\frac{1}{4}, independently of all other throws. Eric throws all three coins at the same time.

Events AA and BB are defined as follows.

AA: all three coins show the same result

BB: at least one of the biased coins shows a head

(a) Show that P(B)=716\mathrm{P}(B)=\frac{7}{16}.

(2)

(b) Find P(A∣B)\mathrm{P}(A\mid B).

(2)

The random variable XX is the number of heads obtained when Eric throws the three coins.

(c) Draw up the probability distribution table for XX.

(3)

题目中文翻译

Eric 有三枚硬币。其中一枚是均匀硬币,另外两枚各自有偏,每次抛掷出现正面的概率均为 14\frac{1}{4},且与所有其他抛掷相互独立。Eric 同时抛掷这三枚硬币。

事件 AA 和 BB 定义如下。

AA:三枚硬币均出现相同的结果

BB:至少一枚有偏硬币出现正面

(a) 证明 P(B)=716\mathrm{P}(B)=\frac{7}{16}。

(2)

(b) 求 P(A∣B)\mathrm{P}(A\mid B)。

(2)

随机变量 XX 表示 Eric 抛掷这三枚硬币时得到的正面数。

(c) 列出 XX 的概率分布表。

(3)

原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

事件B只涉及两枚偏币,恰一正或两正共三条路径,公平币不必分类。

答题过程

展开

For the two biased coins, B occurs on HT, TH or HH. Therefore

P(B)=1434+3414+1414=3+3+116=716.\mathrm{P}(B)=\frac14\frac34+\frac34\frac14+\frac14\frac14 =\frac{3+3+1}{16}=\boxed{\frac7{16}}.

解法二

思路

展开

也可取两枚偏币都反面的补集。

答题过程

展开

The complement of B is that both biased coins show tails. Hence

P(B)=1−(34)2=1−916=716.\mathrm{P}(B)=1-\biggl(\frac34\biggr)^2=1-\frac9{16}=\boxed{\frac7{16}}.

(b)

解法一

思路

展开

A且B只可能三正面,三反面不符合B。

答题过程

展开

Since A∩BA\cap B occurs only when all three coins show heads,

P(A∩B)=121414=132.\mathrm{P}(A\cap B)=\frac12\frac14\frac14=\frac1{32}.

Using (a),

P(A∣B)=1/327/16=114.\mathrm{P}(A\mid B)=\frac{1/32}{7/16}=\boxed{\frac1{14}}.

(c)

解法一

思路

展开

按公平币为正或反分类,偏币正面数服从两次二项分布。

答题过程

展开

Let YY count heads from the two biased coins. Then

P(Y=0)= 916,P(Y=1)= 616,P(Y=2)= 116.\begin{align*} \mathrm{P}(Y=0)=&\,\frac9{16},\\[4mm] \mathrm{P}(Y=1)=&\,\frac6{16},\\[4mm] \mathrm{P}(Y=2)=&\,\frac1{16}. \end{align*}

Conditioning on the fair coin,

P(X=0)=12916=932,\mathrm{P}(X=0)=\frac12\frac9{16}=\frac9{32}, P(X=1)=12916+12616=1532,\mathrm{P}(X=1)=\frac12\frac9{16}+\frac12\frac6{16}=\frac{15}{32}, P(X=2)=12616+12116=732,\mathrm{P}(X=2)=\frac12\frac6{16}+\frac12\frac1{16}=\frac7{32}, P(X=3)=12116=132.\mathrm{P}(X=3)=\frac12\frac1{16}=\frac1{32}.

Thus

xx00112233
P(X=x)\mathrm{P}(X=x)932\frac9{32}1532\frac{15}{32}732\frac7{32}132\frac1{32}